# NaturalBench: Evaluating Vision-Language Models on Natural Adversarial Samples

### v2

@@ Abstract
Vision-language models (VLMs) have made significant progress in recent visual-question-answering (VQA) benchmarks that evaluate complex visio-linguistic reasoning. However, are these models truly effective? In this work, we show that VLMs still struggle with natural images and questions that humans can easily answer, which we term natural adversarial samples. We also find it surprisingly easy to generate these VQA samples from natural image-text corpora using off-the-shelf models like CLIP and ChatGPT. We propose a semi-automated approach to collect a new benchmark, NaturalBench, for reliably evaluating VLMs with 10,000 human-verified VQA samples. Crucially, we adopt a vision-centric design by pairing each question with two images that yield different answers, preventing “blind” solutions from answering without using the images. This makes NaturalBench more challenging than previous benchmarks that can largely be solved with language priors like commonsense knowledge. We evaluate 53 state-of-the-art VLMs on NaturalBench, showing that models like BLIP-3, LLaVA-OneVision, Cambrian-1, InternLM-XC2, Llama3.2-Vision, Molmo, Qwen2-VL, and even the (closed-source) GPT-4o lag 50%-70% behind human performance (which is above 90%). We analyze why NaturalBench is hard from two angles: (1) Compositionality: Solving NaturalBench requires diverse visio-linguistic skills, including understanding attribute bindings, object relationships, and advanced reasoning like logic and counting. To this end, unlike prior work that uses a single tag per sample, we tag each NaturalBench sample with 1 to 8 skill tags for fine-grained evaluation. (2) Biases: NaturalBench exposes severe biases in VLMs, as models often choose the same answer regardless of the image. We show that debiasing can be crucial for VLM performance. Lastly, we apply our benchmark curation method to diverse data sources, including long captions (over 100 words) and non-English languages like Chinese and Hindi, highlighting its potential for dynamic evaluations of VLMs.
@@ 1 Introduction
Recent vision-language models (VLMs) such as GPT-4o [61], GPT-4Vision [60], BLIP-3 (XGen-MM) [79], LLaVA-OneVision [37], InternLM-XC2 [14], Llama3.2-Vision [17], Molmo [10] and Qwen2-VL [76] have markedly improved performance on challenging visual-question-answering (VQA) benchmarks like MMMU [83] and MME [18]. These benchmarks evaluate VLMs across various domains, such as college-level subjects [83, 53], commonsense reasoning [18, 34], diagram comprehension [30, 52], and complex problem-solving in mathematics, coding, physics, and temporal forecasting [54, 18, 50]. Despite their progress, our research identifies a significant gap: these models still struggle with seemingly simple questions about natural images. Figure 1 shows such VQA samples that humans find easy to solve, while even the state-of-the-art models fail. We term these natural adversarial samples [24] for VLMs.
@@ 4 Experimental Results
Table 1: Performance on NaturalBench. We report the performance of 53 leading VLMs on NaturalBench. All models significantly lag behind human performance, with the performance gap (in G-Acc) between humans and models highlighted in red. The latest models, such as BLIP-3 (XGen-MM), Cambrian-1, LLaVA-OneVision, Llama3.2-Vision, Molmo, and Qwen2-VL lag significantly behind humans by 55% to 70%. Even the best closed-source GPT-4o is still 52% behind humans.
@@
Qwen2-VL [76] CLIP-L-14 Qwen2-1.5B 74.1 50.8 54.4 23.4 -68.7
@@ Appendix B NaturalBench Performance
Table 4: Performance on different subsets of NaturalBench. We report the G-Acc performance of 53 leading VLMs on subsets of NaturalBench.
@@
Qwen2-VL [76] CLIP-L-14 Qwen2-1.5B 17.4 22.8 25.6 35.0 23.4
@@ Appendix D Debiasing Analysis
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] The main claims are reflected in the paper’s contributions and scope. See Section 3 to Section 6.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 5.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 3 for detailed benchmark collection pipeline. We have released the code in our project site.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See supplement.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See supplement.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A]
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No] The assets used are all publicly sourced, and therefore explicit consent was not required.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No] The data used are publicly available datasets that do not contain offensive content and with consent for personally identifiable information.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Section 3.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] We pay the participants above the minimum wage with an hourly pay.

### v3

@@ Abstract
Vision-language models (VLMs) have made significant progress in recent visual-question-answering (VQA) benchmarks that evaluate complex visio-linguistic reasoning. However, are these models truly effective? In this work, we show that VLMs still struggle with natural images and questions that humans can easily answer, which we term natural adversarial samples. We also find it surprisingly easy to generate these VQA samples from natural image-text corpora using off-the-shelf models like CLIP and ChatGPT. We propose a semi-automated approach to collect a new benchmark, NaturalBench, for reliably evaluating VLMs with 10,000 human-verified VQA samples. Crucially, we adopt a vision-centric design by pairing each question with two images that yield different answers, preventing “blind” solutions from answering without using the images. This makes NaturalBench more challenging than previous benchmarks that can largely be solved with language priors like commonsense knowledge. We evaluate 55 state-of-the-art VLMs on NaturalBench, showing that models like BLIP-3, LLaVA-OneVision, Cambrian-1, InternLM-XC2, Llama3.2-Vision, Molmo, Qwen2-VL, and even the (closed-source) GPT-4o lag 50%-70% behind human performance (which is above 90%). We analyze why NaturalBench is hard from two angles: (1) Compositionality: Solving NaturalBench requires diverse visio-linguistic skills, including understanding attribute bindings, object relationships, and advanced reasoning like logic and counting. To this end, unlike prior work that uses a single tag per sample, we tag each NaturalBench sample with 1 to 8 skill tags for fine-grained evaluation. (2) Biases: NaturalBench exposes severe biases in VLMs, as models often choose the same answer regardless of the image. We show that debiasing can be crucial for VLM performance. Lastly, we apply our benchmark curation method to diverse data sources, including long captions (over 100 words) and non-English languages like Chinese and Hindi, highlighting its potential for dynamic evaluations of VLMs.
@@ 1 Introduction
Recent vision-language models (VLMs) such as GPT-o3 [63], GPT-4o [61], Gemini-2.5 [62], BLIP-3 (XGen-MM) [81], LLaVA-OneVision [37], InternLM-XC2 [14], Llama3.2-Vision [17], Molmo [10] and Qwen2-VL [78] have markedly improved performance on challenging visual-question-answering (VQA) benchmarks like MMMU [85] and MME [18]. These benchmarks evaluate VLMs across various domains, such as college-level subjects [85, 53], commonsense reasoning [18, 34], diagram comprehension [30, 52], and complex problem-solving in mathematics, coding, physics, and temporal forecasting [54, 18, 50]. Despite their progress, our research identifies a significant gap: these models still struggle with seemingly simple questions about natural images. Figure 1 shows such VQA samples that humans find easy to solve, while even the state-of-the-art models fail. We term these natural adversarial samples [24] for VLMs.
@@ 4 Experimental Results
Table 1: Performance on NaturalBench. We report the performance of 55 leading VLMs on NaturalBench. All models significantly lag behind human performance, with the performance gap (in G-Acc) between humans and models highlighted in red. The latest models, such as BLIP-3 (XGen-MM), Cambrian-1, LLaVA-OneVision, Llama3.2-Vision, Molmo, and Qwen2-VL lag significantly behind humans by 55% to 70%. Even the best closed-source GPT-o3 is still 30% behind humans.
@@
Qwen2-VL [78] NaViT-14 Qwen2-1.5B 74.1 50.8 54.4 23.4 -68.7
@@ Appendix B NaturalBench Performance
Table 4: Performance on different subsets of NaturalBench. We report the G-Acc performance of 55 leading VLMs on subsets of NaturalBench.
@@
Qwen2-VL [78] NaViT-14 Qwen2-1.5B 17.4 22.8 25.6 35.0 23.4